A taxable entity is planning to purchase an equipment that will save 2,000 per year for 5 years. Cost of the equipment is 7,000. For tax purposes, this machine can be depreciated in 3 years using straight line method. Assume tax rate to be 40% and discount rate to be 9%.

a) What is the savings in each of the five years?

b) What is the NPV of the project?

c) Should the entity purchase the equipment?

Please put in Excel and explain
Answer #1

a: Depreciation in each year = Cost of equipment/ Useful life

= 7000/3

= 2333.33

Tax shield on depreciation = Depreciation*40%

= 2333.33 *40%

= 933.33

Annual savings after tax = 5000*(1-0.4) = 3000

Savings in years 1 to 3 = 3000+ 933.33 = 3933.33

Years 4 and 6 = 3000

b: NPV is as follows

Year Cost of equipment Tax shield on depreciation Annual savings Cash flows
0 -7000 -7000
1 933.332 3000 3933.33
2 933.332 3000 3933.33
3 933.332 3000 3933.33
4 3000 3000
5 3000 3000
NPV 7031.49

Since NPV is positive the equipment should be purchased,

WORKINGS

AutoSave Book1 Excel (Product Activation Failed) Sign in File Insert Page Layout Formulas Data Review View Help Tell me what you want to do 수 Share Home Calibri General Conditional Fornat as Cell Insert Delete Fornat FormattingTable Styles- 、. Sort & Find & Editing Paste . 녀 . 의 _ . 로三들經垣臣 Merge & Center-5 . % , : Filter Select Clipboard Alignment Number Styles Cells C2 xCost of equipment Year Cost of equipment Tax shield on depreciation Annual savings Cash flows 0 7000 -SUM(C3:E3) 5000 (1-0.4) SUM(C4:E4) 5000* (1-0.4) SUM(CS:E5) -5000(1-0.4) =SUM(C6:E6) -5000* (1-0.4) SUM(C7:E7) -5000(1-04) =SUM(C8:E8) 2333.33 *40% -2333.33 *4096 -2333.33 *40% 10 NPV -NPV(996,F4f8)-F3 12 13 14 15 16 17 Sheet1 Ready + 100 cENG 10:28 PM